# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

### v1

@@ VI Greedy Home Swapping
Supplementary Figure S15. The impact of the threshold values of the usage of the motorized vehicles. (A) shows the dependence of the average congestion coefficient $I_{t}$ on the threshold values. (B) shows the dependence of the reduction of emission reductions on the threshold values.
X Examples of home swapping in cities other than Shijiazhuang
Home Swapping in Shenzhen. Similar to Shijiazhuang, we first analyzed the spatial distributions of residence and workplace of commuters in Shenzhen’s urban area. Figure S16 shows the heat map of residence and workplace locations in Shenzhen. Among them, Futian, Bao’an, and Luohu Districts are the most bustling areas, with the highest number of residences and workplaces represented by the red areas, while Longgang , Yantian, and Pingshan Districts represented by the blue and green areas on the right have fewer residences and workplaces. By the computing the KL divergence between the two spatial distributions, we found a KL divergence value of 0.114, indicating that the distribution of residences and workplaces is very similar across the city, suggesting a large potential benefit brought by GHS.
Supplementary Figure S16. The spatial distributions of residence and workplace in the urban area in Shenzhen. (A) shows the heat map of residences in Shenzhen; (B) shows the heat map of workplaces in Shenzhen.
Even though the spatial distributions of residence and workplace in Shenzhen is very similar, the city still experiences traffic congestion [10], indicating that long commuting distance of commuters may be a reason. Therefore, we analyzed and showed the distribution of commuting distance of Shenzhen’s commuters in Figure S17. As we can see, the peak commuting distance for most commuters exceeds 10 km, with more than half of the commuters experiencing long-distance commutes daily, suggesting a significant room for exploitation. It is worth noting that the commuting distances used here are the distances obtained by navigation devices.
Supplementary Figure S17. The distribution of commuting distance in Shenzhen.
Due to the limitation of the Shenzhen’s data without detailed individual trajectory information, it is impossible to infer household membership through individual trajectory analysis. Therefore, we constructed single and double households with a ratio of 6:4 based on statistics from the Shenzhen 2015 City Population Census Yearbook. Initially, commuters living at the same latitude and longitude in the data were randomly paired to form double-commuter families, from which 2806394 individuals were randomly selected to form double-commuter households, leaving the remaining commuters as single-commuter households. Results show that $\Delta d$ after GHS reached 62%.
Supplementary Figure S18. $\Delta d$ obtained after considering single socio-demographic factor in Greedy Home Swapping (GHS) in Shenzhen.
Next, we used Shenzhen’s data on amenity (i.e. point of interest (POI)), housing price, and commuter routes based on the output from navigation devices to analyze the socio-demographic GHS, as shown in Figure S18. From Figures S18A and B, it can be observed that after c-GHS and p-GHS, $\Delta d$ improves as the restrictions are relaxed. However, changes in the dissimilarity in amenity accessibility in a-GHS did not significantly enhance $\Delta d$ in Figure S18C, likely due to data granularity, with data spatial resolution at the road level and POIs calculated around road surroundings, leading to indistinguishable road in some cases. In addition, as we see in Figure S18B, there is a notable increase in $\Delta d$ when the tolerance on the difference of housing price increases from 10% to 30%, which aligns with the wide range of housing prices in Shenzhen.
We also examine the joint consideration of all the socio-demographic factors in GHS in Shenzhen similar to the case of Shijiazhuang in the main text. As shown in Figure S19A, under low tolerance ($T_{center}=1km$, $T_{price}=10\%$, $T_{amenity}=0.1$), $\Delta d$ is less than 10%. As tolerance increase, $\Delta d$ noticeably improves, consistent with the conclusions in Figure 4D of the main text. However, under high tolerance ($T_{center}=5km$, $T_{price}=50\%$, $T_{amenity}=0.5$), $\Delta d$ in Shenzhen exceeds 40%, significantly higher than that obtained in Shijiazhuang. Such high reduction in commuting distance of over 40% may be due to the large disparity in housing price in Shenzhen, creating a barriers for households to choose the most convenient residence locations, leading to an extensive commuting by a large set of commuters. The result of adding new activity centers in the Shenzhen simulation is consistent with the results of Shijiazhuang, as shown in Figure S19B.
Supplementary Figure S19. $\Delta d$ of Shenzhen after SDGHS. (A) The reduction in commuting distance after SDGHS with a different tolerances. (B) The reduction in commuting distance after SDGHS with a medium tolerance and a different number of new activity centers.
Home Swapping in Abidjan. We utilize data from Abidjan, the most economically prosperous center in Cote d’Ivoire, for the simulation of home swapping. First, based on the location of commuters’ residence and workplace, we visualize the spatial distributions of residence and workplace in Figure S20, where the transition from orange to deep red indicates areas with a higher concentration of residences and workplaces, suggesting a high population density. Conversely, green and blue areas represent locations with fewer residences and workplaces. By computing the KL divergence between the two spatial distributions, we found that the KL divergence is 0.113, suggesting distributions of residence and workplace in Abidjan are highly similar.
Supplementary Figure S20. The spatial distributions of residence and workplace. (A) shows the heat map of residence distribution in Abidjan; (B) shows the heat map of workplace distribution in Abidjan.
The distributions of residence and workplace in Abidjan are similar, implying a significant room for exploitation using GHS. The distribution of individual commuting distance of commuters in Abidjan is shown in Figure S21. As we can see, most commuters travel 5 km to their work, and there are a large number of commuters who travel for a commuting distance exceeds 10 km.
Supplementary Figure S21. The distribution of commuting distance in Abidjan.
After simulation of GHS in Abidjan, we found that single-commuter household achieved an 83.66% reduction in commuting distance. Since we cannot infer household membership from this simple data, we used the same methodto form double-commuter householdsby randomly pairing individuals living at the same latitude and longitude, resulting in a reduction of 41.36% of commuting distance after GHS as shown in Figure S22A.
Supplementary Figure S22. $\Delta d$ after GHS. (A) shows the reduction in commuting distance after GHS for households of different size; (B) shows $\Delta d$ as a function of the ratio of single-commuter to double-commuter households.
Next, we examine the reduction in commuting distance as a function of households of different size in Figure S22A. As we can see, $\Delta d$ significantly decreases from single-commuter to double-commuter households, while it decreases slowly for households with three, four, or five commuters. The highest $\Delta d$ is found in single-commuter households, and both single and double-commuter households can represent the overall benefit brought by GHS to the city. It is because the city is not composed purely of single or double-commuter families, but rather a certain proportion of single and double-commuter households. Figure S22B shows $\Delta d$ after GHS as a function of the ratio of single-commuter to double-commuter households, showing that as the proportion of single-commuter households increases, $\Delta d$ gradually increases. This suggests that, in comparison to small- or medium-sized cities, metropolitan areas with a higher percentage of single-commuter households have a higher potential for reducing commuting distance by GHS.
Supplementary Figure S23. $\Delta d$ of Abidjan after c-GHS. (A) $\Delta d$ after c-GHS. (B) $\Delta d$ after c-GHS with $T_{center}=3km$ for different number of new activity centers.
From Figure S23A, it can be observed that after c-GHS, $\Delta d$ improves as the restrictions are relaxed. Then we examined the effect of adding new activity centers, which are established by the most-visited locations as new activity centers at a distance of 3km from the original center. In Figure S23B, the result after c-GHS with new activity centers in Abidjan is consistent with the results of Shijiazhuang in the main text.

### v2

@@ X Impact of Commuting Mode Shift Threshold
XI The Impact of Induced Trips after Home Swapping
Any traffic optimization leads to the emergence of new induced trips [12, 13, 14, 15]. Therefore, we conducted preliminary analyses to examine how induced trips might impact our results. Using 15-minute discrete mobility trajectories, we identified potential induced travelers by analyzing both commuters who shifted their departure times (8108-10 AM $\rightarrow$ 787-8 AM) and non-commuters who adjusted their travel schedules. We simulated how congestion relief might generate new trips by randomly selecting $5\%-25\%$ of individuals who moved more than $2.5$ km and classifying them as motorized travelers, proportional to actual commuter volumes, and incorporating their real trajectory data. Each synthetic traveler was assigned a morning departure time (7-8 AM) matching actual commuting patterns, with routes generated from their recorded stay points through navigation data. Through 10 repeated sampling iterations to ensure robustness, we found that while induced travel does gradually increase average congestion coefficients (Figure S17A), the values remain substantially lower than real-world baselines. The reduction in carbon emissions decreases to $71.7\%$ when considering $25\%$ induced trips (Figure S17B), though the system maintains significant benefits since average commuting distances (11.1km) far exceed induced trip lengths (7.6km). These results demonstrate that even accounting for temporal shifts in travel behavior, our home-swapping approach delivers meaningful improvements. We recognize that our current data cannot fully capture potential spatial induced trips (completely new routes enabled by reduced congestion), though the strong periodicity of weekday travel patterns suggests such effects would be limited. This comprehensive analysis confirms the robustness of our findings while transparently acknowledging the boundaries of our current modeling framework.
Supplementary Figure S17. Effects of induced morning travel on urban congestion and emission reduction. (A) City-wide average congestion coefficients under different induced travel proportions during 7-8 AM, where the induced travel proportion is defined as the ratio of newly added travelers to the original number of commuters. (B) Corresponding reduction in carbon emissions under varying induced travel proportions during the same morning peak period. All results represent averages from 10 independent simulation runs.
XII Impact of Multi-modal Transportation on Home Swapping
We have addressed the over-simplification of assuming a single commuting mode by expanding our analysis to incorporate non-motorized transport, buses, and private cars, guided by empirical data from the 2017 Beijing Transport Development Annual Report [16]. Nevertheless, we remarked that the major city studied in the main text, i.e. Shijiazhuang, did not have a metro system at the time the dataset was obtained, and commuters were classified into three modes: trips under $2.5$ km as non-motorized, trips over $2.5$ km with an average speed $\leq 3$ m/s (consistent with reported bus speeds) as bus commuters, and trips over $2.5$ km with an average speed >3>3 m/s as private car users. This categorization aligned closely with official statistics ($18.5\%$ bus, $45\%$ car vs. reported $22.3\%$ bus, $37\%$ car), ensuring validity.
Based on the above analysis, the commuting mode of each commuter can be determined. If a commuter still travels a distance greater than $2.5$ km after the home swapping process, it is assumed that the commuter still maintains the original commuting mode. To model multi-modal impacts, we integrated the city’s entire bus network - $246$ routes, $3661$ stops, and real-time schedules - into our simulations. This paper demonstrates that all bus routes within the urban area of Shijiazhuang were characterized using real-world stop trajectories and departure time data (Figure S18A). Bus trajectories during morning peak hours were reconstructed using navigation data, while car traffic flows were combined with bus traffic flows to compute congestion and emissions. It is important to note that the average length of a bus is four times that of a vehicle. Furthermore, given that the bus uses diesel, the carbon emission factor is 0.281 $kgCO_{2}e/km$.
Supplementary Figure S18. (A) Spatial distribution of all 246 bus routes (3,661 stops) in the Shijiazhuang, with route density indicated by color intensity (darker colors represent higher route density). Actual bus trajectories and schedules were incorporated from city transit data (see detail in SI section 1). (B) The reduction of average commuting distance by home-swapping (GHS and SDGHS) under multi-modal commuting conditions (medium tolerance). (C) The reduction of carbon emissions during morning peak hours under multi-modal conditions, with both GHS and SDGHS showing significant decreases in total city-wide carbon emissions despite accounting for bus and private vehicle interactions.
Results in Figure S18B demonstrate that the core findings, with the assumption of multi-modal transportation, remain robust: GHS reduced the average commuting distances by  $50\%$, while SDGHS (with medium tolerance for heterogeneity) achieved a $13\%$ reduction. Morning peak congestion decreased from 0.61 (empirical baseline) to $0.35$ (GHS) and $0.38$ (SDGHS), with corresponding declines in carbon emissions, $83\%$ (GHS) and $43.5\%$ (SDGHS), similar to that under the single-mode transportation (Figure S18C). These outcomes confirm that our proposed home-swapping approaches retain efficacy across diverse transportation modes.
XIII Commuting Surplus Analysis
On the other hand, it is useful to incorporate a consumer surplus [17] analysis into our framework to quantify the economic impacts of home-swapping. We first define the fuel cost $T_{i}$ of an individual $i$ traveling through a set of paths $Path_{i}(t)$ is given by
$T_{i}=P_{gos}\times{\textstyle\sum_{\alpha\in Path_{i}(t)}d_{i}^{\alpha}\times f_{i}^{\alpha}}$ (18)
where $P_{gos}$ represents the fuel price ($7$ CNY per liter); $Path_{i}(t)$ is the set of road segments traversed by individual $i$ during time interval $t$; $d_{i}^{\alpha}$ is the length of road segment $\alpha$; $f_{i}^{\alpha}$ stands for the fuel consumption rate (liters per kilometer) of individual $i$ on road segment $\alpha$. Fuel consumption rates were parameterized according to roadway congestion levels: $0.165$ L/km (severely congested), $0.125$ L/km (moderately congested), and $0.085$ L/km (free-flow) [18], as derived from real-time traffic coefficient data.
We then define the gain in daily commuting of individual $i$ after home-swapping to be
$T_{i}^{SL}=T_{i}^{Real}-T_{i}^{SDGHS}$ (19)
which is the difference between $T_{i}^{Real}$, the real daily commuting cost for individual $i$, and $T_{i}^{SDGHS}$, the daily commuting cost for individual $i$ after home-swapping under the SDGHS framework.
The gain $T_{i}^{SL}$ - defined as the difference between the commuting cost before and after home-swapping - reveals that over $80\%$ of households gain economic benefits under medium tolerance conditions (Figure S19), with monthly savings distributed broadly across income groups. This aligns with urban economics theory, as swaps can benefit individual households, even given our constraints to minimize housing cost disruptions (i.e. price differences $15\%$) and to ensure most households retain comparable housing utility.
Supplementary Figure S19. Distribution of the commuting surplus $T_{i}^{SL}$ defined as the difference between original and post-swap commuting costs following the SDGHS optimization, with benefits sustained across income levels under medium ($15\%$) housing price change constraints. Over $80\%$ of households achieved a positive commuting surplus, demonstrating widespread economic benefits from the residential swapping process.
XIV Examples of Home Swapping in the China’s Major Cities
Similar to Shijiazhuang, the initial analysis of the distributions of residence and workplace of commuters was conducted in the China’s 28 major cities (see detail in Section 1). According to ref. [6], the individual’s residence is designated as the trip origin, and workplace is designated as the trip destination. The KL divergence between these distributions ranged from 0.06 to 0.18, with an average of 0.087 (Figure S20A), indicating that the distribution of residence and workplace is very similar across cities. Figure S20B presents heatmaps for Lanzhou, Nanjing, and Beijing. Similar to the pattern observed in Shijiazhuang, both residential and workplace locations are concentrated near the city center, while suburban areas show the opposite trend.
Supplementary Figure S20. Residential and Workplace Distribution in the China’s 28 major cities. (A) KL divergence scatter plot of residential and workplace distribution. (B) The spatial distributions of residential and workplace in Lanzhou, Nanjing, and Beijing.
Due to the limitation of the origin-destination (OD) data without detailed individual trajectory information, it is impossible to infer household membership. Therefore, we constructed a realistic composite of 1-commuter and 2-commuter households based on the proportions reported in the 2019 China Statistical Yearbook [19], as shown in Supplementary Table S2. Initially, commuters living at the same grid cell in the data were randomly paired to form 2-commuter households. A proportional random sample of these was subsequently extracted, with the remaining commuters as 1-commuter households. A decline in average commuting distance was observed after GHS across the 28 major cities of varying sizes (ranking by Seventh National Population Census [20]). Notably, the correlation between the percentage decrease in average commuting distance ($\Delta d$) and city population size appears slightly positive. Interestingly, Chongqing showed an unusually high percentage decrease ($\Delta d=88\%$) attributable to its unique mountainous topography, which suggests limitation on the applicability and accuracy of conventional planar geographic information systems for generating navigation. Due to the similar limitation of POI data and housing price data, we excluded Chongqing from the analysis of GHS with socio-demographic factors.
Supplementary Figure S21. GHS with Socio-demographic Consideration. (A) The percentage decrease in average commuting distance after c-GHS, p-GHS, and a-GHS in the China’s 28 major cities. (B) The percentage decrease in average commuting distance after SDGHS and the impact of adding new centers in China’s top-10 congested cities. All results represent averages from 10 independent simulations per city
Then, we used data on amenities (POI), housing prices, and navigation routes of the 28 cities to analyze GHS considering a single socio-demographic. Given the $1$ km grid resolution, the constraints on distance to the city center, housing prices, and amenity accessibility were relaxed for household relocation. In addition to reducing commuting distance, two households swapped their homes only if: (1) The difference in their distance to city center (defined as the most visited grid cell between 7-8 AM) is less than 3 km (c-GHS); (2) The difference between their housing prices is less than $30\%$ (p-GHS); (3) The difference in amenity accessibility between the two residences is less than 0.3 (a-GHS). Amenity accessibility was calculated based solely on seven types of amenities within each grid cell, without considering neighboring cells. As shown in Figure S21A, the observed percentage decrease in average commuting distance after c-GHS, p-GHS, and a-GHS are detailed for the China’s 28 major cities, which are ordered by population size.
As shown in Figure S21B, we also examine the SDGHS for China’s top-10 congested cities, based on data from the Baidu Map Group’s China Urban Transportation Report [21]. In this analysis, Shijiazhuang replaces Chongqing due to the latter’s special mountainous topography, as it is invalid to describe the regional features by using POI and housing price. To further investigate the impact of adding secondary centers, we identified the highest-traffic points in each administrative district and selected the top four locations outside the primary center as new centers. The results show that five-center cities achieved a greater reduction in average commuting distances after SDGHS than their single-center counterparts, consistent with the findings from Shijiazhuang.
Finally, we applied the same methodology to estimate the carbon emission reduction after GHS for the 10 most congested cities of China. Since the resolution of OD data is $1$ hour, it lacks the accurate individual departure time from his/her dwelling. To address this, we inferred departure time based on the OD matrix for the 7-8 AM period and actual commuting route navigation data. Specifically, if the navigated commute time exceeded one hour, the individual was assumed to depart at 7 AM. If it was less than one hour, a departure time was randomly assigned within the feasible window to ensure arrival by 8 AM. Notably, when considering socio-demographic factors, the results of OD data show a greater improvement than those obtained by the high-resolution trajectories of Shijiazhuang in the main text. It can be attributed to several factors: Firstly, the coarse granularity of the data, constrained by a $1$ km $\times$ 11 km grid resolution, homogenizes POI and housing price data within the grid, thereby reducing household heterogeneity to loosen the restrictions of home swapping. Secondly, due to the limited temporal resolution of the OD data, the individual departure time is estimated, not the real record, which underestimates the temporal crowdedness to flatten the peak-hour congestion. Thirdly, the current household matching mechanism is much more simplified and does not fully account for the complexity of real-world household structures, which lowered the difficulty of home swapping. Finally, the temporal coverage of the OD data may fail to capture cross-hour commuters outside the observation window.
 City #\#grids #\#population (million)  Ratio ($\frac{\#2-\text{commuter}}{\#1-\text{commuter}}$)
Shanghai 2442 24.871 1.44
Beijing 8717 21.893 1.37
Chengdu 5609 20.938 1.36
Guangzhou 3801 18.677 0.82
Shenzhen 1455 17.56 0.82
Tianjin 5671 13.866 1.99
Xi’an 2991 12.953 1.66
Zhengzhou 5507 12.601 1.82
Wuhan 4353 12.327 1.65
Hangzhou 5910 11.936 1.33
Shijiazhuang 3825 11.235 2.39
Changsha 5054 10.048 1.71
Harbin 2499 10.001 2.22
Hefei 1115 9.369 1.79
Nanjing 3880 9.315 1.82
Jinan 1052 9.202 2.14
Shenyang 1789 9.07 1.94
Changchun 2020 9.065 2.13
Nanning 5674 8.742 1.19
Kunming 3492 8.46 1.45
Fuzhou 7862 8.291 1.23
Nanchang 1733 6.255 2.03
Guiyang 1909 5.987 1.46
Taiyuan 1847 5.304 2.05
Lanzhou 2322 4.359 1.85
Hohhot 6152 3.446 2.39
Haikou 610 2.873 1.46
Lhasa 903 0.868 1.32
Supplementary Table S2. The attributes of the China’s 28 major cities.